what is x and y-intercept for each problem
x+3y=9


and 2x+3y=6




and 4x-5y=10


and x +y =5

Answer by rfer(16322)

y int, x=0
x int, y=0
------------
x+3y=9
0+3y=9
y=3
(0,3)
-------------
x+3*0=9
x=9
(9,0)
you try the others.